OpenStudy (anonymous):

what is the simplified form of 4x^2-25/2x+5

OpenStudy (jim766):

note the numerator is a difference of 2 squares

OpenStudy (callisto):

Factorize the numerator using the identity \(a^2 -b^2 = (a+b)(a-b)\), then cancel the common factor.

OpenStudy (jim766):

if you know what that is, makes it easier to factor

OpenStudy (anonymous):

okk. then ?

OpenStudy (jim766):

then cancel out what is 4x^2 - 25 factored

OpenStudy (anonymous):

(2x-5)(2x+5).. ?

OpenStudy (jim766):

right, now (2x + 5)(2x -5) / 2x + 5 what is left after you cancel out 2x + 5

OpenStudy (anonymous):

(2x-5) . ii know that . but how do i find the restrictions.?

OpenStudy (jim766):

keep in mind the fractions are undefined when the denominator = 0 so we need to ask ourself what would make 2x + 5 = 0 , just solve it

OpenStudy (anonymous):

2x+5=0 2x=-5 x= -5/2

OpenStudy (jim766):

right, the only time that this cant be done is when x = -5/2 or -2.5 make sense?

OpenStudy (anonymous):

ohh okk ii get it. thankyou. youve been a great help !!!!! (:

OpenStudy (jim766):

yw